OpenStudy (anonymous):

An elevator named L in "63 Building" in Seoul is notorious for its low speed. It takes 3 seconds for L just to move one floor. It takes another 20 seconds for it to have its door open and close automatically. What is even more frustrating, is that the elevator travels only to odd-numbered floors, and the doors will automatically open and close at all odd numbered floors. How long will it take (in seconds) for L leaving the 1st floor "just to arrive" at the 15th floor? Do not include the time taken for the doors to open / close at the 1st or 15th floor.

OpenStudy (anonymous):

Well it takes 3 seconds for it to move from one floor to another. So it takes 3 * 14 = 42 seconds for it just to go from the first to 15th floor. But on the way it will take 20 seconds to open and close on each odd numbered floors. Since it's "leaving" the first floor I assume that the first floor is not accounted for the 20 second delay. Similarly, we don't count the 20 sec delay on the 15th floor either because the question says "just to arrive". So basically the 20 sec delay occurs on the following floors: 3,5,7,9,11,13 which is 20*6=120. Hence the time taken for it leave the first floor and "just arrive" the 15th floor is 42 + 120 = 162 seconds.

OpenStudy (anonymous):

Lol I didn't notice the "Do not include the time taken for the doors to open / close at the 1st or 15th floor." until after answering but I still did exactly that :D

OpenStudy (anonymous):

you too :D

OpenStudy (anonymous):

i am puting another easy question @genius12